# ALINE: Joint Amortization for Bayesian Inference and Active Data Acquisition

### v1

@@ 1 Introduction
Bayesian inference [26] and Bayesian experimental design [58] offer principled mathematical means for reasoning under uncertainty and for strategically gathering data, respectively. While both are foundational, they both introduce notorious computational challenges. For example, in scenarios with continuous data streams, repeatedly applying gold-standard inference methods such as Markov Chain Monte Carlo (MCMC) [11] to update posterior distributions can be computationally demanding, leading to various approximate sequential inference techniques [8, 16], yet challenges in achieving both speed and accuracy persist. Similarly, in Bayesian experimental design (BED) or Bayesian active learning (BAL), the iterative estimation and optimization of design objectives can become costly, especially in sequential learning tasks requiring rapid design decisions [19, 28, 52], such as in psychophysical experiments where the goal is to quickly infer the subject’s perceptual or cognitive parameters [66, 54]. Moreover, a common approach in BED involves greedily optimizing for single-step objectives, such as the Expected Information Gain (EIG), which measures the anticipated reduction in uncertainty. However, this leads to myopic designs that may be suboptimal, not fully considering how current choices influence future learning opportunities [21].
@@
In this paper, we introduce Amortized Active Learning and INference Engine (Aline), a novel framework designed to overcome these limitations by unifying amortized Bayesian inference and active data acquisition within a single, cohesive system (Table 1). Aline utilizes a transformer-based architecture [65] that, in a single forward pass, concurrently performs posterior estimation, generates posterior predictive distributions, and decides which data point to query next. Critically, and in contrast to existing methods, Aline offers flexible, targeted acquisition: it can dynamically adjust at runtime its data-gathering strategy to focus on any specified combination of model parameters or predictive tasks. This is enabled by an attention mechanism allowing the policy to condition on specific inference goals, making it particularly effective in the presence of nuisance variables [55] or for focused investigations. Aline is trained using a self-guided reinforcement learning objective; the reward is the improvement in the log-probability of its own approximate posterior over the selected targets, a principle derived from variational bounds on the expected information gain [22]. Extensive experiments on diverse tasks demonstrate Aline’s ability to simultaneously deliver fast, accurate inference and rapidly propose informative data points.
@@ 2 Background
Consider a parametric conditional model defined on some space $\mathcal{Y}\subseteq\mathbb{R}^{d_{\mathcal{Y}}}$ of output variables $y$ given inputs (or covariates) $x\in\mathcal{X}\subseteq\mathbb{R}^{d_{\mathcal{X}}}$, and parameterized by $\theta\in\Theta\subseteq\mathbb{R}^{L}$. Let $\mathcal{D}_{T}=\{(x_{i},y_{i})\}_{i=1}^{T}$ be a collection of $T$ data points (or context) and $p(\mathcal{D}_{T}\,|\,\theta)\equiv p(y_{1:T}\,|\,x_{1:T},\theta)$ denote the likelihood function associated with the model. Given a prior distribution $p(\theta)$, the classical Bayesian inference or prediction problem involves estimating either the posterior distribution $p(\theta\,|\,\mathcal{D}_{T})\propto p(\mathcal{D}_{T}\,|\,\theta)p(\theta)$, or the posterior predictive distribution $p(y_{1:M}^{\star}\,|\,x_{1:M}^{\star},\mathcal{D}_{T})=\mathbb{E}_{p(\theta\,|\,\mathcal{D}_{T})}[p(y_{1:M}^{\star}\,|\,x_{1:M}^{\star},\theta,\mathcal{D}_{T})]$ over target outputs ${y}^{\star}_{1:M}:=({y}^{\star}_{1},...,{y}^{\star}_{M})$ corresponding to a given set of target inputs $x^{\star}_{1:M}:=(x^{\star}_{1},...,x^{\star}_{M})$. Estimating these quantities repeatedly via approximate inference methods such as MCMC can be computationally costly [26], motivating the need for amortized inference methods.
@@
BED [44, 12, 60, 58] methods aim to sequentially select the next input (or design parameter) $x$ to query, in order to maximize the Expected Information Gain (EIG), that is the information gained about parameters $\theta$ upon observing $y$:
@@ 3 Amortized active learning and inference engine
To achieve both instant, informative querying and accurate inference, we propose to jointly learn an amortized inference model $q_{\phi}$ and an acquisition policy $\pi_{\psi}$ within a single, integrated architecture. Given the accumulated data history $\mathcal{D}_{t-1}$ and a specific target $\xi\in\Xi$, the policy $\pi_{\psi}$ selects the next query $x_{t}$ designed to be most informative for that target. Subsequently, the new data point $(x_{t},y_{t})$ is observed, and the inference model $q_{\phi}$ updates its estimate of the corresponding posterior or posterior predictive distribution. A conceptual workflow of Aline is illustrated in Figure 1. In the remainder of this section, we detail the objectives for training the inference network $q_{\phi}$ (Section 3.1) and the acquisition policy $\pi_{\psi}$ (Section 3.2), discuss their practical implementation (Section 3.3), and describe the unified model architecture (Section 3.4).
@@ 3.1 Amortized inference
We use the inference network $q_{\phi}$ to provide accurate approximations of the true Bayesian posterior $p(\theta\,|\,\mathcal{D}_{T})$ or posterior predictive distributions $p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$, given the acquired data $\mathcal{D}_{T}$. We train $q_{\phi}$ via maximum-likelihood (Eq. 1). Specifically, for parameter targets $\xi=\xi_{S}^{\theta}$, our objective is:
$\mathcal{L}_{S}^{\theta}(\phi)=-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)}\left[\log q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})\right]=-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)}\left[\sum_{l\in S}\log q_{\phi}(\theta_{l}\,|\,\mathcal{D}_{T})\right],$ (5)
where we adopt a diagonal or mean field approximation, where the joint distribution is obtained as a product of marginals $q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})=\prod_{l\in S}q_{\phi}(\theta_{l}\,|\,\mathcal{D}_{T})$. Analogously, for predictive targets $\xi=\xi_{p_{\star}}^{{y}^{\star}}$, we assume a factorized likelihood over targets sampled from the target input distribution $p_{\star}(x^{\star})$:
$\mathcal{L}^{{y}^{\star}}_{p_{\star}}(\phi)=-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)p_{\star}(x^{\star}_{1:M})p({y}^{\star}_{1:M}\,|\,x^{\star}_{1:M},\theta)}\left[\sum_{m=1}^{M}\log q_{\phi}({y}^{\star}_{m}\,|\,x^{\star}_{m},\mathcal{D}_{T})\right].$ (6)
The factorized form of these training objectives is a common scalable choice in the neural process literature [24, 49, 50, 13] and more flexible than it might seem, as conditional marginal distributions can be extended to represent full joints autoregressively [50, 9, 13]. For simplicity and tractability, within the scope of this paper we focus on the marginals, leaving the autoregressive extension to future work. Eqs. 5 and 6 form the basis for training the inference component $q_{\phi}$. Optimizing them minimizes the Kullback-Leibler (KL) divergence between the true target distributions (posterior or predictive) defined by the underlying generative process and the model’s approximations $q_{\phi}$ [49]. Learning an accurate $q_{\phi}$ is crucial as it not only determines the quality of the final inference output but also serves as the basis for guiding the data acquisition policy, as we see next.
@@ 3.2 Amortized data acquisition
For completeness, we include a derivation of $\text{sEIG}_{\theta_{S}}$ in Section A.1 which is analogous to that of sEIG in [21]. Directly optimizing sEIG is generally intractable due to its reliance on the unknown true posterior $p(\theta_{S}\,|\,\mathcal{D}_{T})$. We circumvent this by substituting $p(\theta_{S}\,|\,\mathcal{D}_{T})$ with its approximation $q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})$ (from Eq. 5), yielding the tractable objective $\mathcal{J}^{\theta}_{S}$ for training $\pi_{\psi}$:
@@
$\text{sEPIG}(\psi)$ $:=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}\,|\,\pi_{\psi})}[H[p({y}^{\star}\,|\,x^{\star})]-H[p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})]]$
@@
This result adapts Theorem 1 in [21] for the predictive case (see Section A.2 for proof) and, unlike single-step EPIG (Eq. 4), considers the entire trajectory $\mathcal{D}_{T}$ given the policy $\pi_{\psi}$.
Now, similar to Eq. 8, we use the inference network $q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$ to replace the true posterior predictive distribution $p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$ in Proposition 1 to obtain our active learning objective:
@@
Finally, the following proposition proves that our acquisition objectives, $\mathcal{J}_{S}^{\theta}$ and $\mathcal{J}_{p_{\star}}^{{y}^{\star}}$, are variational lower bounds on the true total information gains ($\text{sEIG}_{\theta_{S}}$ and sEPIG, respectively), making them principled tractable objectives for our goal. The proof is given in Section A.3.
@@
This principle of using approximate posterior (or predictive) distributions to bound information gain is foundational in Bayesian experimental design (e.g., [22]) and has been extended to sequential amortized settings (e.g., the sCEE bound in [7]). Maximizing these $\mathcal{J}$ objectives thus encourages policies that increase information about the targets. The tightness of these bounds is governed by the expected KL divergence between the true quantities and their approximation, with a more accurate $q_{\phi}$ leading to tighter bounds and a more effective training signal for the policy.
@@ 3.3 Training methodology for Aline
which maximizes the expected cumulative $\gamma$-discounted reward over trajectories [64].
@@
For the per-step rewards $R_{t}$ to be meaningful, the inference network $q_{\phi}$ must provide accurate estimates of posteriors or predictive distributions at each intermediate step $t$ of the acquisition sequence, not just at the final step $T$. Consequently, the practical training objective for $q_{\phi}$, denoted $\mathcal{L}_{\text{NLL}}(\phi)$, encourages this step-wise accuracy. In practice, training proceeds in episodes. For each episode: (1) A ground truth parameter set $\theta$ is sampled from the prior $p(\theta)$. (2) A target specifier $\xi$ is sampled from $p(\xi)$. (3) If the target is predictive ($\xi=\xi_{p_{*}}^{{y}^{\star}}$), $M$ target input locations $\{x^{\star}_{m}\}_{m=1}^{M}$ are sampled from $p_{\star}(x^{\star})$, and their corresponding true outcomes $\{{y}^{\star}_{m}\}_{m=1}^{M}$ are simulated using $\theta$, similarly as [63]. The negative log-likelihood loss $\mathcal{L}_{\text{NLL}}(\phi)$ for $q_{\phi}$ in an episode is then computed by averaging over the $T$ acquisition steps and predictions, using the Monte Carlo estimates of the objectives defined in Eqs. 5 and 6:
$\mathcal{L}_{\text{NLL}}(\phi)\approx\begin{cases}-\frac{1}{T}\frac{1}{|S|}\sum_{t=1}^{T}\sum_{l\in S}\log q(\theta_{l}\,|\,\mathcal{D}_{t}),&\text{if }\xi=\xi^{\theta}_{S}\\ -\frac{1}{T}\frac{1}{M}\sum_{t=1}^{T}\sum_{m=1}^{M}\log q({y}^{\star}_{m}\,|\,x^{\star}_{m},\mathcal{D}_{t}),&\text{if }\xi=\xi^{{y}^{\star}}_{p_{\star}}.\\ \end{cases}$ (12)
@@
To ensure $q_{\phi}$ provides a reasonable reward signal early in training, we employ an initial warm-up phase where only $q_{\phi}$ is trained, with data acquisition $(x_{t},y_{t})$ guided by random actions instead of $\pi_{\psi}$. After the warm-up, $q_{\phi}$ and $\pi_{\psi}$ are trained jointly. A detailed step-by-step training algorithm is provided in Section B.1.

### v2

@@ 1 Introduction
Bayesian inference [28] and Bayesian experimental design [61] offer principled mathematical means for reasoning under uncertainty and for strategically gathering data, respectively. While both are foundational, they introduce notorious computational challenges. For example, in scenarios with continuous data streams, repeatedly applying gold-standard inference methods such as Markov Chain Monte Carlo (MCMC) [13] to update posterior distributions can be computationally demanding, leading to various approximate sequential inference techniques [10, 18], yet challenges in achieving both speed and accuracy persist. Similarly, in Bayesian experimental design (BED) or Bayesian active learning (BAL), the iterative estimation and optimization of design objectives can become costly, especially in sequential learning tasks requiring rapid design decisions [21, 30, 55], such as in psychophysical experiments where the goal is to quickly infer the subject’s perceptual or cognitive parameters [70, 57]. Moreover, a common approach in BED involves greedily optimizing for single-step objectives, such as the Expected Information Gain (EIG), which measures the anticipated reduction in uncertainty. However, this leads to myopic designs that may be suboptimal, which do not fully consider how current choices influence future learning opportunities [23].
@@
In this paper, we introduce Amortized Active Learning and INference Engine (Aline), a novel framework designed to overcome these limitations by unifying amortized Bayesian inference and active data acquisition within a single, cohesive system (Table˜1). Aline utilizes a transformer-based architecture [69] that, in a single forward pass, concurrently performs posterior estimation, generates posterior predictive distributions, and decides which data point to query next. Critically, and in contrast to existing methods, Aline offers flexible, targeted acquisition: it can dynamically adjust at runtime its data-gathering strategy to focus on any specified combination of model parameters or predictive tasks. This is enabled by an attention mechanism allowing the policy to condition on specific inference goals, making it particularly effective in the presence of nuisance variables [58] or for focused investigations. Aline is trained using a self-guided reinforcement learning objective; the reward is the improvement in the log-probability of its own approximate posterior over the selected targets, a principle derived from variational bounds on the expected information gain [24]. Extensive experiments on diverse tasks demonstrate Aline’s ability to simultaneously deliver fast, accurate inference and rapidly propose informative data points.
@@ 2 Background
Consider a parametric conditional model defined on some space $\mathcal{Y}\subseteq\mathbb{R}^{d_{\mathcal{Y}}}$ of output variables $y$ given inputs (or covariates) $x\in\mathcal{X}\subseteq\mathbb{R}^{d_{\mathcal{X}}}$, and parameterized by $\theta\in\Theta\subseteq\mathbb{R}^{L}$. Let $\mathcal{D}_{T}=\{(x_{i},y_{i})\}_{i=1}^{T}$ be a collection of $T$ data points (or context) and $p(\mathcal{D}_{T}\,|\,\theta)\equiv p(y_{1:T}\,|\,x_{1:T},\theta)$ denote the likelihood function associated with the model, which we assume to be well-specified in this paper (i.e., it matches the true data generation process). Given a prior distribution $p(\theta)$, the classical Bayesian inference or prediction problem involves estimating either the posterior distribution $p(\theta\,|\,\mathcal{D}_{T})\propto p(\mathcal{D}_{T}\,|\,\theta)p(\theta)$, or the posterior predictive distribution $p(y_{1:M}^{\star}\,|\,x_{1:M}^{\star},\mathcal{D}_{T})=\mathbb{E}_{p(\theta\,|\,\mathcal{D}_{T})}[p(y_{1:M}^{\star}\,|\,x_{1:M}^{\star},\theta,\mathcal{D}_{T})]$ over target outputs ${y}^{\star}_{1:M}:=({y}^{\star}_{1},...,{y}^{\star}_{M})$ corresponding to a given set of target inputs $x^{\star}_{1:M}:=(x^{\star}_{1},...,x^{\star}_{M})$. Estimating these quantities repeatedly via approximate inference methods such as MCMC can be computationally costly [28], motivating the need for amortized inference methods.
@@
BED [47, 14, 63, 61] methods aim to sequentially select the next input (or design parameter) $x$ to query in order to maximize the Expected Information Gain (EIG), that is, the information gained about parameters $\theta$ upon observing $y$:
@@
Related work.
A major family of ABI methods is Neural Processes (NPs) [27, 26], that learn a mapping from the observed context data points to a predictive distribution for new target points. Early NPs often employed MLP-based encoders [27, 26, 38, 11], while recent works utilize more advanced attention and transformer architectures [42, 53, 52, 19, 20, 5, 4]. Complementary to NPs, methods within simulation-based inference [17] focus on amortizing the posterior distribution [54, 49, 34, 60, 51, 75]. More recently, methods for amortizing both the posterior and posterior predictive distributions have been proposed [31, 59, 15]. Specifically, ACE [15] shows how to flexibly condition on diverse user-specified inference targets, a method Aline incorporates for its own flexible inference capabilities. Building on this principle of goal-directed adaptability, Aline advances it by integrating a learned policy that dynamically tailors the data acquisition strategy to the specified objectives. Existing amortized BED or BAL methods [23, 40, 9, 46] that learn an offline design policy do not provide real-time estimates of the posterior, unlike Aline. Recent exceptions include methods like RL-sCEE [8] and vsOED [65], that use a variational posterior bound of EIG to provide amortized posterior inference via a separate proposal network. Compared to these methods, Aline uses a single, unified architecture where the same model performs amortized inference for both posterior and posterior predictive distributions, and learns the flexible acquisition policy.
@@ 3 Amortized active learning and inference engine
To achieve both instant, informative querying and accurate inference, we propose to jointly learn an amortized inference model $q_{\phi}$ and an acquisition policy $\pi_{\psi}$ within a single, integrated architecture. Given the accumulated data history $\mathcal{D}_{t-1}$ and a specific target $\xi\in\Xi$, the policy $\pi_{\psi}$ selects the next query $x_{t}$ designed to be most informative for that target. Subsequently, the new data point $(x_{t},y_{t})$ is observed, and the inference model $q_{\phi}$ updates its estimate of the corresponding posterior or posterior predictive distribution. A conceptual workflow of Aline is illustrated in Figure˜1. In the remainder of this section, we detail the objectives for training the inference network $q_{\phi}$ (Section˜3.1) and the acquisition policy $\pi_{\psi}$ (Section˜3.2), discuss their practical implementation (Section˜3.3), and describe the unified model architecture (Section˜3.4).
@@ 3.1 Amortized inference
We use the inference network $q_{\phi}$ to provide accurate approximations of the true Bayesian posterior $p(\theta\,|\,\mathcal{D}_{T})$ or posterior predictive distribution $p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$, given the acquired data $\mathcal{D}_{T}$. We train $q_{\phi}$ via maximum-likelihood (Eq. 1). Specifically, for parameter targets $\xi=\xi_{S}^{\theta}$, our objective is:
$\mathcal{L}_{S}^{\theta}(\phi)=-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)}\left[\log q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})\right]\approx-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)}\left[\sum_{l\in S}\log q_{\phi}(\theta_{l}\,|\,\mathcal{D}_{T})\right],$ (5)
where we adopt a diagonal or mean field approximation, where the joint distribution is obtained as a product of marginals $q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})\approx\prod_{l\in S}q_{\phi}(\theta_{l}\,|\,\mathcal{D}_{T})$. Analogously, for predictive targets $\xi=\xi_{p_{\star}}^{{y}^{\star}}$, we assume a factorized likelihood over targets sampled from the target input distribution $p_{\star}(x^{\star})$:
$\mathcal{L}^{{y}^{\star}}_{p_{\star}}(\phi)$ $=-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)p_{\star}(x^{\star}_{1:M})p({y}^{\star}_{1:M}\,|\,x^{\star}_{1:M},\theta)}\left[\log q_{\phi}({y}^{\star}_{1:M}\,|\,x^{\star}_{1:M},\mathcal{D}_{T})\right]$
$\approx-\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\theta)p_{\star}(x^{\star}_{1:M})p({y}^{\star}_{1:M}\,|\,x^{\star}_{1:M},\theta)}\left[\sum_{m=1}^{M}\log q_{\phi}({y}^{\star}_{m}\,|\,x^{\star}_{m},\mathcal{D}_{T})\right].$ (6)
The factorized form of these training objectives is a common scalable choice in the neural process literature [26, 52, 53, 15] and is more flexible than it might seem, as conditional marginal distributions can be extended to represent full joints autoregressively [53, 11, 15]. However, a full autoregressive model would require multiple forward passes to compute the reward signal for our policy at each training step, making the learning process computationally intractable. Therefore, for simplicity and tractability, within the scope of this paper we focus on the marginals, leaving the autoregressive extension to future work. Eqs. 5 and 3.1 form the basis for training the inference component $q_{\phi}$. Optimizing them minimizes the Kullback-Leibler (KL) divergence between the true target distributions (posterior or predictive) defined by the underlying generative process and the model’s approximations $q_{\phi}$ [52]. Learning an accurate $q_{\phi}$ is crucial as it not only determines the quality of the final inference output but also serves as the basis for guiding the data acquisition policy, as we see next.
@@ 3.2 Amortized data acquisition
For completeness, we include a derivation of $\text{sEIG}_{\theta_{S}}$ in Section˜A.1 which is analogous to that of sEIG in [23]. Directly optimizing sEIG is generally intractable due to its reliance on the unknown true posterior $p(\theta_{S}\,|\,\mathcal{D}_{T})$. We circumvent this by substituting $p(\theta_{S}\,|\,\mathcal{D}_{T})$ with its approximation $q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})$ (from Eq. 5), yielding the tractable objective $\mathcal{J}^{\theta}_{S}$ for training $\pi_{\psi}$:
@@
$\text{sEPIG}(\psi)$ $:=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}_{T}\,|\,\pi_{\psi})}[H[p({y}^{\star}\,|\,x^{\star})]-H[p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})]]$
@@
This result adapts Theorem 1 in [23] for the predictive case (see Section˜A.2 for proof) and, unlike single-step EPIG (Eq. 4), considers the entire trajectory $\mathcal{D}_{T}$ given the policy $\pi_{\psi}$.
Now, similar to Eq. 8, we use the inference network $q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$ to replace the true posterior predictive distribution $p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$ in Proposition˜1 to obtain our active learning objective:
@@
Finally, the following proposition proves that our acquisition objectives, $\mathcal{J}_{S}^{\theta}$ and $\mathcal{J}_{p_{\star}}^{{y}^{\star}}$, are variational lower bounds on the true total information gains ($\text{sEIG}_{\theta_{S}}$ and sEPIG, respectively), making them principled tractable objectives for our goal. The proof is given in Section˜A.3.
@@
This principle of using approximate posterior (or predictive) distributions to bound information gain is foundational in Bayesian experimental design (e.g., [24]) and has been extended to sequential amortized settings [8, 65]. Maximizing these $\mathcal{J}$ objectives thus encourages policies that increase information about the targets. The tightness of these bounds is governed by the expected KL divergence between the true quantities and their approximation, with a more accurate $q_{\phi}$ leading to tighter bounds and a more effective training signal for the policy. Additionally, our objectives solely rely on Aline’s variational posterior, which does not require an explicit likelihood, making it naturally applicable to problems with implicit likelihoods where only forward sampling is possible.
@@ 3.3 Training methodology for Aline
which maximizes the expected cumulative $\gamma$-discounted reward over trajectories [68]. Gradients from this policy loss only update the policy parameters $\psi$. They are not propagated back to the inference network $q_{\phi}$ to ensure each component has a clear and distinct objective.
@@
For the per-step rewards $R_{t}$ to be meaningful, the inference network $q_{\phi}$ must provide accurate estimates of posteriors or predictive distributions at each intermediate step $t$ of the acquisition sequence, not just at the final step $T$. Consequently, the practical training objective for $q_{\phi}$, denoted $\mathcal{L}_{\text{NLL}}(\phi)$, encourages this step-wise accuracy. In practice, training proceeds in episodes. For each episode: (1) A ground truth parameter set $\theta$ is sampled from the prior $p(\theta)$. (2) A target specifier $\xi$ is sampled from $p(\xi)$. (3) If the target is predictive ($\xi=\xi_{p_{\star}}^{{y}^{\star}}$), $M$ target inputs $\{x^{\star}_{m}\}_{m=1}^{M}$ are sampled from $p_{\star}(x^{\star})$, and their corresponding true outcomes $\{{y}^{\star}_{m}\}_{m=1}^{M}$ are simulated using $\theta$, similarly as [67]. The negative log-likelihood loss $\mathcal{L}_{\text{NLL}}(\phi)$ for $q_{\phi}$ in an episode is then computed by averaging over the $T$ acquisition steps and predictions, using the Monte Carlo estimates of the objectives defined in Eqs. 5 and 3.1:
[…]
@@
To ensure $q_{\phi}$ provides a reasonable reward signal early in training, we employ an initial warm-up phase where only $q_{\phi}$ is trained, with data acquisition $(x_{t},y_{t})$ guided by random actions instead of $\pi_{\psi}$. After the warm-up, $q_{\phi}$ and $\pi_{\psi}$ are trained jointly. A detailed step-by-step training algorithm is provided in Section˜B.1.
Figure 2: The Aline architecture. The model takes historical observations (context set), candidate points (query set), and the current inference goal (target set) as inputs. These are transformed by embedding layers and subsequently by transformer layers. Finally, an acquisition head determines the next data point to query, while an inference head performs the approximate Bayesian inference.